OpenStudy (shelld93):

1/2 + 4/9 I'm getting 5/18 but my calculator says 17/18 and I don't know where I'm going wrong.

OpenStudy (anonymous):

Convert to a common denominator. Try 18

OpenStudy (shelld93):

yeah ive got 18 as my denominator but my numerator is wrong according to my calculator and i dont understand where im going wrong :/

OpenStudy (anonymous):

\[\huge \frac{ 1 }{ 2 } = \frac{ ? }{ 18 }\]

OpenStudy (shelld93):

17? 16? i really do not know.

OpenStudy (shelld93):

the online course is pulling numbers out of thin air ..

OpenStudy (anonymous):

17/18

OpenStudy (anonymous):

it's not thin air rofl. \[\huge \frac{ 1 }{ 2 } = \frac{ 1*9 }{ 2*9 } = \frac{ ? }{ 18 }\]

OpenStudy (shelld93):

is that times ?

OpenStudy (shelld93):

freaking seems it.

OpenStudy (anonymous):

it looks like it mine is plus

OpenStudy (anonymous):

\[is \it \]

OpenStudy (shelld93):

okay hang on. I need to show working so from 1/2 + 4/9 gives the common denominator of 18

OpenStudy (shelld93):

now what do I do to get the 17 that you got ?

OpenStudy (anonymous):

yeah the *s are times

OpenStudy (anonymous):

I'm sure aspen has it though

OpenStudy (anonymous):

9 times what = 18

OpenStudy (shelld93):

2

OpenStudy (anonymous):

4* 2 =?

OpenStudy (shelld93):

8

OpenStudy (anonymous):

2 times 18 = what

OpenStudy (shelld93):

36

OpenStudy (anonymous):

2times what = 18

OpenStudy (shelld93):

9

OpenStudy (anonymous):

9+8+?

OpenStudy (anonymous):

re-read this if you still don't get it http://www.mathsisfun.com/fractions_addition.html

OpenStudy (anonymous):

9+8=?

OpenStudy (shelld93):

17

OpenStudy (shelld93):

thanks you !!!!!

OpenStudy (anonymous):

your welcome